A lottery has one $2000. prize, two $1000. prizes, and ten $500. prizes. Two thousand tickets are sold at $5.00 each. find the expectation if a person buys two tickets. Four choices are -$.50, -$1.00, -$2.00, -$.25
How do I even go about wotking this?
Ellen
: A lottery has one $2000. prize, two $1000. prizes, and ten $500. prizes. Two thousand tickets are sold at $5.00 each. find the expectation if a person buys two tickets. Four choices are -$.50, -$1.00, -$2.00, -$.25
How do I even go about wotking this?
Ellen

Answer by stanbon(17981)

A lottery has one $2000. prize, two $1000. prizes, and ten $500. prizes. Two thousand tickets are sold at $5.00 each. find the expectation if a person buys two tickets. Four choices are -$.50, -$1.00, -$2.00, -$.25
-------------------
The random variable values are 2000, 1000, 500, -5
The corresponding probabilities are (2/2000),(4/2000),(20/2000),(1976/2000)
---------------
The expected value is (2/2000)*2000 + (4/2000)*1000 + (20/2000)*500 + (1976/2000)*(-5) = 2 + 2 + 5 -4.94 = $4.06
=================
Cheers,
Stan H.